Find the standard form of the hyperbola with verticies (0,3) and (0,-3) and asmptotes y=3/2x and y=-3/2x
Answer by lwsshak3(11628)

Find the standard form of the hyperbola with verticies (0,3) and (0,-3) and asymptotes y=3/2x and y=-3/2x
**
This is a hyperbola with vertical transverse axis.
Its standard form of equation: (y-k)^2/a^2-(x-h)^2/b^2=1 , (h for given hyperbola:
center: (0,0)
length of vertical transverse axis=6=2a
a=3
a^2=9
..
slope of asymptotes=a/b=3/2
b=2a/3=2
b^2=4
..
Equation of given hyperbola:
y^2/9-x^2/4=1